OpenStudy (anonymous):

can someone please solve 3(x+1) +5 = 3x+2

OpenStudy (anonymous):

3x+3+5 = 3x+2 8=2 No solutions

OpenStudy (anonymous):

3x+3+5=3x+2 0x=-6 x=error